Show all the solutions. Round off answers to two decimal places.

3. High-voltage power lines are familiar sights throughout the country. The aluminum wire used for some of these lines has a diameter of 2.5 cm. What is the resistance of 10,000 m of this wire? Resistivity = 2.83 x 10-8 ohm-m.



1
Expert's answer
2021-04-22T07:29:14-0400

Radius of wire, "r=\dfrac{2.5}{2}=1.25\space cm=1.25\times10^{-2} m"

Area of cross section, "A" = "\pi r^2=4.90\times10^{-4}\space m^2"

Length of wire, "l=10,000\space m"

Resistivity, "\rho=2.83\times 10^{-8}\space ohm-m"

Resistance, "R=\rho\dfrac{l}{A}=2.83\times10^{-8}\times\dfrac{10000}{4.90\times10^{-4}}"

"R=5.77\space \Omega"